In an odd twist of events, I lack a ruler but do have a pair of wires, a power supply,
and an ammeter. I know that if I run a current through the wires I can balance the
downward force of gravity on the top wire with the repulsive magnetic force between
them. I know the weight per unit length of the wires is 0.15 N/m. I can use the power
supply and the ammeter to choose the current (assumed to be the same current in
each wire). What value of current should I choose to make the distance between the
two wires equal to exactly 1.00 cm? Note that these are parallel wires and not a single
wire in an external magnetic field.
1
Expert's answer
2017-03-28T05:39:06-0400
The force between two wires can be divided as F = 2m0*I1*I2/(4*3.14*r), or I = sqrt(2*3.14*r*F/(m0)) = 87 A